# A contramodule generalization of Neeman’s flat and projective module theorem

### v1

@@ 3. Topological Rings
The definition of a contramodule over a topological ring goes back to [19, Remark A.3]. We refer to [22, Section 2.1], [28, Sections 1.2 and 5], and [26, Sections 7–8] for further details and references in connection with the definitions and assertions presented below in this section. The pedagogical exposition in [29, Sections 6–7] is particularly recommended.
@@
Now let $\mathfrak{R}$ be a complete, separated topological ring with a base of neighborhoods of zero formed by open right ideals. Then the functor $\mathfrak{R}[[{-}]]\colon\mathsf{Sets}\longrightarrow\mathsf{Sets}$ has a natural structure of a monad on the category of sets (in the sense of [17, Chapter VI]). For any set $X$, the monad unit $\epsilon_{\mathfrak{R},X}\colon X\longrightarrow\mathfrak{R}[[{-}]]$ is the map taking every element $x\in X$ to the formal linear combination $\sum_{y\in X}r_{y}y$ with $r_{x}=1$ and $r_{y}=0$ for $y\neq x$. The monad multiplication $\phi_{\mathfrak{R},X}\colon\mathfrak{R}[[\mathfrak{R}[[X]]]]\longrightarrow\mathfrak{R}[[X]]$ is the “opening of parentheses” map assigning a formal linear combination to a formal linear combination of formal linear combinations. In order to open the parentheses, one needs to compute pairwise products of elements of the ring $\mathfrak{R}$ and certain infinite direct sums in $\mathfrak{R}$, which are understood as the limits of finite partial sums in the topology of $\mathfrak{R}$. The conditions imposed on $\mathfrak{R}$ (including in particular the requirement of a topology base of open right ideals) guarantee the convergence.
In the general category-theoretic terminology, one usually speaks about “algebras over monads”, but in our context we prefer to call them modules over monads. A left $\mathfrak{R}$-contramodule is defined as a module over the monad $X\longmapsto\mathfrak{R}[[X]]$ on the category of sets. In other words, a left $\mathfrak{R}$-contramodule $\mathfrak{P}$ is a set endowed with a left contraaction map $\pi_{\mathfrak{P}}\colon\mathfrak{R}[[\mathfrak{P}]]\longrightarrow\mathfrak{P}$ satisfying the following contraassociativity and contraunitality axioms. Firstly, the map $\mathfrak{R}[[\pi_{\mathfrak{P}}]]\colon\mathfrak{R}[[\mathfrak{R}[[\mathfrak{P}]]\longrightarrow\mathfrak{R}[[\mathfrak{P}]]$ induced by the map $\pi_{\mathfrak{P}}$ and the “opening of parentheses” map $\phi_{\mathfrak{R},\mathfrak{P}}\colon\mathfrak{R}[[\mathfrak{R}[[\mathfrak{P}]]\longrightarrow\mathfrak{R}[[\mathfrak{P}]]$ must have equal compositions with the contraaction map $\pi_{\mathfrak{P}}$,
@@
The contratensor product $\mathcal{N}\odot_{\mathfrak{R}}\mathfrak{P}$ of a discrete right $\mathfrak{R}$-module $\mathcal{N}$ and left $\mathfrak{R}$-contramodule $\mathfrak{P}$ is an abelian group constructed as the cokernel of (the difference of) the natural pair of maps
@@
A left $\mathfrak{R}$-contramodule $\mathfrak{F}$ is called flat if the contratensor product functor ${-}\odot_{\mathfrak{R}}\mathfrak{F}\colon{\operatorname{\mathsf{Discr--}}}\mathfrak{R}\longrightarrow\mathsf{Ab}$ is exact. It follows that the class of all flat $\mathfrak{R}$-contramodules is preserved by the directed colimits in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ (because the directed colimits are exact in $\mathsf{Ab}$), and all the projective $\mathfrak{R}$-contramodules are flat. We will denote the full subcategory of flat left $\mathfrak{R}$-contramodules by $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}\subset\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$.
@@ 4. Countably Presentable Contramodules
The exposition in this section is based on [28, Section 6] and [26, Sections 8–10]. In the case of a topological ring with a base of neighborhoods of zero formed by two-sided ideals, [21, Section D.1] is a relevant reference.
@@
(a) all left $\mathfrak{R}$-contramodules are complete (but they need not be separated);
@@
Part (a) is [26, Lemma 8.4(b)]. Part (b) is [26, Lemma 8.8]. ∎
@@
Let $\mathfrak{R}$ be a complete, separated topological ring, and let $\mathfrak{R}\supset\mathfrak{I}_{1}\supset\mathfrak{I}_{2}\supset\mathfrak{I}_{3}\supset\dotsb$ be a countable base of neighborhoods of zero in $\mathfrak{R}$ consisting of open two-sided ideals $\mathfrak{I}_{n}$,  $n\geq 1$. Denote by $R_{n}=\mathfrak{R}/\mathfrak{I}_{n}$ the corresponding discrete quotient rings. Then the rule assigning to a flat left $\mathfrak{R}$-contramodule $\mathfrak{F}$ the collection of $R_{n}$-modules $F_{n}=\mathfrak{F}/(\mathfrak{I}_{n}\rightthreetimes\mathfrak{F})$ is an equivalence between the category of flat left $\mathfrak{R}$-contramodules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ and the category formed by the following sets of data:
@@
This is [26, Proposition 9.1(a)]. ∎
@@
This is [26, Lemmas 9.4 and 9.5, and Theorem 10.1]. ∎
@@
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Then any countably presentable flat $\mathfrak{R}$-contramodule has projective dimension at most $1$ in the abelian category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$.
@@ 5. Acyclic Complexes of Flat Contramodules are Contraacyclic
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open right ideals. We will say that a complex of flat left $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is pure acyclic if it is acyclic in the exact category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$[26, Section 11]. Equivalently, a complex of flat $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is pure acyclic if and only if $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is acyclic in the abelian category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ and the $\mathfrak{R}$-contramodules of cocycles of $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ are flat.
@@
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Then, for any complex of projective left $\mathfrak{R}$-contramodules $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ and any pure acyclic complex of flat left $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, any morphism of complexes $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero. In other words, any pure acyclic complex of flat $\mathfrak{R}$-contramodules is contraacyclic (as a complex in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$).
@@
Notice, first of all, that any projective object in $\mathsf{B}$ is a direct summand of a coproduct of copies of $P$; hence any complex of projective objects in $\mathsf{B}$ is a direct summand of a complex whose components are coproducts of copies of $P$. For the purposes of this paper, we are really interested in the case $\kappa=\aleph_{1}$. When $\kappa=\aleph_{0}$, the category $\mathsf{B}$ is the module category $\mathsf{B}\simeq R{\operatorname{\mathsf{--Mod}}}$, where $R$ is (the opposite ring to) the endomorphism ring of the small projective generator $P\in\mathsf{B}$. This case is covered by [33, Proposition 4.3] (cf. [18, Construction 4.3]). In the case of an uncountable regular cardinal $\kappa$, the desired assertion was established in [30, second proof of Proposition 7.2]. ∎
@@
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Then the category of pure acyclic complexes of flat left $\mathfrak{R}$-contramodules is $\aleph_{1}$-accessible. The $\aleph_{1}$-presentable objects of this category are precisely all the pure acyclic complexes of countably presentable flat left $\mathfrak{R}$-contramodules.
@@ 6. Contraacyclic Complexes of Flat Contramodules are Acyclic
A left $\mathfrak{R}$-contramodule $\mathfrak{C}$ is said to be cotorsion [28, Definition 7.3] if one has $\operatorname{Ext}^{\mathfrak{R},1}(\mathfrak{F},\mathfrak{C})=0$ for every flat left $\mathfrak{R}$-contramodule $\mathfrak{F}$. The following lemma may help the reader to feel more comfortable.
@@
This is a part of [28, Corollary 7.8] (see the terminology in [28, Section 3]). ∎
@@
In view of Lemma 6.4, we have a short exact sequence of complexes of abelian groups $0\longrightarrow\operatorname{Hom}^{\mathfrak{R}}(\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}^{\mathfrak{R}}(\mathfrak{Q}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}^{\mathfrak{R}}(\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow 0$. So the dual argument to the proof of Lemma 2.4(a) applies. ∎
@@
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open right ideals. Then every complex of flat left $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be included into a short exact sequence $0\longrightarrow\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow 0$ in $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$ with a contractible complex of flat cotorsion $\mathfrak{R}$-contramodules $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ and a complex of flat $\mathfrak{R}$-contramodules $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$.
@@
Let $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be the direct sum of contractible two-terms complexes $0\longrightarrow\overline{\mathfrak{C}}^{n}\overset{{\mathrm{id}}}{\longrightarrow}\overline{\mathfrak{C}}^{n}\longrightarrow 0$ with the cohomological grading defined by the rule that $\mathfrak{C}^{n}=\overline{\mathfrak{C}}^{n}\oplus\overline{\mathfrak{C}}^{n+1}$. Let $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be the morphism of complexes whose components are the monomorphisms $\mathfrak{F}^{n}\longrightarrow\overline{\mathfrak{C}}^{n}$ and the compositions $\mathfrak{F}^{n}\longrightarrow\mathfrak{F}^{n+1}\longrightarrow\overline{\mathfrak{C}}^{n+1}$. Then $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is a (termwise) monomorphism of complexes of $\mathfrak{R}$-contramodules. Let $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be the cokernel complex. Then there are short exact sequences of $\mathfrak{R}$-contramodules $0\longrightarrow\overline{\mathfrak{C}}^{n+1}\longrightarrow\mathfrak{G}^{n}\longrightarrow\overline{\mathfrak{G}}^{n}\longrightarrow 0$. Once again, Lemma 4.2 tells us that the $\mathfrak{R}$-contramodules $\mathfrak{G}^{n}$ are flat as extensions of flat $\mathfrak{R}$-contramodules. ∎

### v3

@@ 3. Topological Rings
The definition of a contramodule over a topological ring goes back to [27, Remark A.3]. We refer to [30, Section 2.1], [38, Sections 1.2 and 5], and [34, Sections 7–8] for further details and references in connection with the definitions and assertions presented below in this section. The pedagogical exposition in [39, Sections 6–7] is particularly recommended.
@@
Now let $\mathfrak{R}$ be a complete, separated topological ring with a base of neighborhoods of zero formed by open right ideals. Then the functor $\mathfrak{R}[[{-}]]\colon\mathsf{Sets}\longrightarrow\mathsf{Sets}$ has a natural structure of a monad on the category of sets (in the sense of [25, Chapter VI]). For any set $X$, the monad unit $\epsilon_{\mathfrak{R},X}\colon X\longrightarrow\mathfrak{R}[[X]]$ is the map taking every element $x\in X$ to the formal linear combination $\sum_{y\in X}r_{y}y$ with $r_{x}=1$ and $r_{y}=0$ for $y\neq x$. The monad multiplication $\phi_{\mathfrak{R},X}\colon\mathfrak{R}[[\mathfrak{R}[[X]]]]\longrightarrow\mathfrak{R}[[X]]$ is the “opening of parentheses” map assigning a formal linear combination to a formal linear combination of formal linear combinations. In order to open the parentheses, one needs to compute pairwise products of elements of the ring $\mathfrak{R}$ and certain infinite direct sums in $\mathfrak{R}$, which are understood as the limits of finite partial sums in the topology of $\mathfrak{R}$. The conditions imposed on $\mathfrak{R}$ (including, in particular, the requirement of a topology base of open right ideals) guarantee the convergence.
In the general category-theoretic terminology, one usually speaks about “algebras over monads”, but in our context we prefer to call them modules over monads. A left $\mathfrak{R}$-contramodule is defined as a module over the monad $X\longmapsto\mathfrak{R}[[X]]$ on the category of sets. In other words, a left $\mathfrak{R}$-contramodule $\mathfrak{P}$ is a set endowed with a left contraaction map $\pi_{\mathfrak{P}}\colon\mathfrak{R}[[\mathfrak{P}]]\longrightarrow\mathfrak{P}$ satisfying the following contraassociativity and contraunitality axioms. Firstly, the map $\mathfrak{R}[[\pi_{\mathfrak{P}}]]\colon\mathfrak{R}[[\mathfrak{R}[[\mathfrak{P}]]]]\longrightarrow\mathfrak{R}[[\mathfrak{P}]]$ induced by the map $\pi_{\mathfrak{P}}$ and the “opening of parentheses” map $\phi_{\mathfrak{R},\mathfrak{P}}\colon\mathfrak{R}[[\mathfrak{R}[[\mathfrak{P}]]]]\longrightarrow\mathfrak{R}[[\mathfrak{P}]]$ must have equal compositions with the contraaction map $\pi_{\mathfrak{P}}$,
@@
The contratensor product $\mathcal{N}\odot_{\mathfrak{R}}\mathfrak{P}$ of a discrete right $\mathfrak{R}$-module $\mathcal{N}$ and a left $\mathfrak{R}$-contramodule $\mathfrak{P}$ is an abelian group constructed as the cokernel of (the difference of) the natural pair of maps
@@
A left $\mathfrak{R}$-contramodule $\mathfrak{F}$ is called flat if the contratensor product functor ${-}\odot_{\mathfrak{R}}\nobreak\mathfrak{F}\colon\allowbreak{\operatorname{\mathsf{Discr--}}}\mathfrak{R}\longrightarrow\mathsf{Ab}$ is exact. It follows that the class of all flat $\mathfrak{R}$-contramodules is preserved by the directed colimits in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ (because the directed colimits are exact in $\mathsf{Ab}$), and all the projective $\mathfrak{R}$-contramodules are flat. We will denote the full subcategory of flat left $\mathfrak{R}$-contramodules by $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}\subset\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$.
@@ 4. Countably Presentable Contramodules
The exposition in this section is based on [38, Section 6] and [34, Sections 8–10]. In the case of a topological ring with a base of neighborhoods of zero formed by two-sided ideals, [29, Section E.1] is a relevant reference.
@@
(a) all left $\mathfrak{R}$-contramodules are complete (but not necessarily separated);
@@
Part (a) is [34, Lemma 8.4(b)]. Part (b) is [34, Lemma 8.8]. ∎
@@
Let $\mathfrak{R}$ be a complete, separated topological ring, and let $\mathfrak{R}\supset\mathfrak{I}_{1}\supset\mathfrak{I}_{2}\supset\mathfrak{I}_{3}\supset\dotsb$ be a countable base of neighborhoods of zero in $\mathfrak{R}$ consisting of open two-sided ideals $\mathfrak{I}_{n}$,  $n\geq 1$. Denote by $R_{n}=\mathfrak{R}/\mathfrak{I}_{n}$ the corresponding discrete quotient rings. Then the rule assigning to a flat left $\mathfrak{R}$-contramodule $\mathfrak{F}$ the collection of $R_{n}$-modules $F_{n}=\mathfrak{F}/(\mathfrak{I}_{n}\rightthreetimes\mathfrak{F})$ is an equivalence between the category of flat left $\mathfrak{R}$-contramodules $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$ and the category of following sets of data:
@@
This is [34, Proposition 9.1(a)]. ∎
Assuming that the topological ring $\mathfrak{R}$ has a countable base of neighborhoods of zero consisting of open right ideals, we will say that a left $\mathfrak{R}$-contramodule $\mathfrak{P}$ is countably generated if $\mathfrak{P}$ is a quotient contramodule of a free contramodule $\mathfrak{R}[[X]]$ spanned by a countable set $X$. This is equivalent to $\mathfrak{P}$ being an $\aleph_{1}$-generated object of $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ in the sense of [1, Definition 1.67].
@@
This is [34, Lemmas 9.4 and 9.5, and Theorem 10.1]. ∎
@@
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Then any countably presentable flat $\mathfrak{R}$-contramodule has projective dimension $\leq\nobreak 1$ in the abelian category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$.
@@ 5. Acyclic Complexes of Flat Contramodules are Contraacyclic
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open right ideals. We will say that a complex of flat left $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is pure acyclic if it is acyclic in the exact category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$[34, Section 11]. Equivalently, a complex of flat $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is pure acyclic if and only if $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is acyclic in the abelian category $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$ and the $\mathfrak{R}$-contramodules of cocycles of $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ are flat. Let us denote the full subcategory of pure acyclic complexes of flat left $\mathfrak{R}$-contramodules by $\mathsf{Ac}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})\subset\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})\subset\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$.
@@
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Then, for any complex of projective left $\mathfrak{R}$-contramodules $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ and any pure acyclic complex of flat left $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, any morphism of complexes $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is homotopic to zero. In other words, any pure acyclic complex of flat $\mathfrak{R}$-contramodules is contraacyclic (as a complex in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$, or equivalently, in $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}}$).
@@
Notice, first of all, that any projective object in $\mathsf{B}$ is a direct summand of a coproduct of copies of $P$; hence any complex of projective objects in $\mathsf{B}$ is a direct summand of a complex whose components are coproducts of copies of $P$. For the purposes of this paper, we are really interested in the case $\kappa=\aleph_{1}$. When $\kappa=\aleph_{0}$, the category $\mathsf{B}$ is the module category $\mathsf{B}\simeq R{\operatorname{\mathsf{--Mod}}}$, where $R$ is (the opposite ring to) the endomorphism ring of the small projective generator $P\in\mathsf{B}$. This case is covered by [44, Proposition 4.3] (cf. [26, Construction 4.3]). In the case of an uncountable regular cardinal $\kappa$, the desired assertion was established in [40, second proof of Proposition 7.2]. ∎
@@
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open two-sided ideals. Then the category $\mathsf{Ac}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ of pure acyclic complexes of flat left $\mathfrak{R}$-contramodules is $\aleph_{1}$-accessible. A pure acyclic complex of flat $\mathfrak{R}$-contramodules $\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is $\aleph_{1}$-presentable as an object of $\mathsf{Ac}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}}_{\mathsf{flat}})$ if and only if all the $\mathfrak{R}$-contramodules $\mathfrak{H}^{n}$,  $n\in\mathbb{Z}$, are countably presentable.
@@
Remark 5.4.
Theorem 5.1 is a contramodule generalization of the similar assertion in Neeman’s paper [26, Theorem 8.6], which is the module version. Our proof of Theorem 5.1 is inspired by, but independent of the proof of the related assertion in [26].
Alternatively, it is possible to deduce our Theorem 5.1 from Neeman’s result. To this end, one argues as follows. Let $\mathfrak{R}\supset\mathfrak{I}_{1}\supset\mathfrak{I}_{2}\supset\mathfrak{I}_{3}\supset\dotsb$ be a countable base of neighborhoods of zero in $\mathfrak{R}$ consisting of open two-sided ideals $\mathfrak{I}_{n}$,  $n\geq 1$ (as in Lemma 4.4 and Proposition 4.6). Put $R_{n}=\mathfrak{R}/\mathfrak{I}_{n}$.
Then, by Lemma 4.4, for any flat left $\mathfrak{R}$-contramodules $\mathfrak{P}$ and $\mathfrak{F}$, there is a natural isomorphism of abelian groups
[…]
where the transition maps in the projective system are induced by the functors $R_{n}\otimes_{R_{n+1}}{-}\,\colon R_{n+1}{\operatorname{\mathsf{--Mod}}}\longrightarrow R_{n}{\operatorname{\mathsf{--Mod}}}$. When the $\mathfrak{R}$-contramodule $\mathfrak{P}$ is projective, the $R_{n}$-modules $\mathfrak{P}/\mathfrak{I}_{n}\rightthreetimes\mathfrak{P}$ are projective, too, and it follows that any $R_{n}$-module map $\mathfrak{P}/\mathfrak{I}_{n}\rightthreetimes\mathfrak{P}\longrightarrow\mathfrak{F}/\mathfrak{I}_{n}\rightthreetimes\mathfrak{F}$ can be lifted to an $R_{n+1}$-module map $\mathfrak{P}/\mathfrak{I}_{n+1}\rightthreetimes\mathfrak{P}\longrightarrow\mathfrak{F}/\mathfrak{I}_{n+1}\rightthreetimes\mathfrak{F}$. So the transition maps in the projective system are surjective.
Now, for any complex of projective left $\mathfrak{R}$-contramodules $\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ and any complex of flat left $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$, we have a natural isomorphism of complexes of abelian groups
[…]
where the transition maps are surjective in every degree. Whenever the complex of flat $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is pure acyclic, so are the complexes of flat $R_{n}$-modules $\mathfrak{F}/\mathfrak{I}_{n}\rightthreetimes\mathfrak{F}$ (by Lemma 4.3(a)), so the complexes $\operatorname{Hom}_{R_{n}}^{\text{\smaller\smaller$\scriptstyle\bullet$}}(\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}}/\mathfrak{I}_{n}\rightthreetimes\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\>\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}/\mathfrak{I}_{n}\rightthreetimes\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ are acyclic by [26, Theorem 8.6]. It remains to point out that the projective limit of a sequence (indexed by the positive integers) of termwise surjective maps of acyclic complexes of abelian groups is an acyclic complex again. Thus the complex $\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is acyclic whenever the complex $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is pure acyclic.
@@ 6. Contraacyclic Complexes of Flat Contramodules are Acyclic
A left $\mathfrak{R}$-contramodule $\mathfrak{C}$ is said to be cotorsion [38, Definition 7.3] if one has $\operatorname{Ext}^{\mathfrak{R},1}(\mathfrak{F},\mathfrak{C})=0$ for every flat left $\mathfrak{R}$-contramodule $\mathfrak{F}$. We denote the full subcategory of cotorsion left $\mathfrak{R}$-contramodules by $\mathfrak{R}{\operatorname{\mathsf{--Contra}}}^{\mathsf{cot}}\subset\mathfrak{R}{\operatorname{\mathsf{--Contra}}}$. The following lemma may help the reader to feel more comfortable.
@@
This is a part of [38, Corollary 7.8] (see the terminology in [38, Section 3] or in Section 7 below). ∎
@@
In view of Lemma 6.4, we have a short exact sequence of complexes of abelian groups $0\longrightarrow\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{H}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{Q}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{P}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}})\longrightarrow 0$. So the dual argument to the proof of Lemma 2.5(a) applies, proving that the complex of abelian groups $\operatorname{Hom}^{\mathfrak{R},{\text{\smaller\smaller$\scriptstyle\bullet$}}}(\mathfrak{T}^{\text{\smaller\smaller$\scriptstyle\bullet$}},\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}})$ is acyclic. ∎
@@
Let $\mathfrak{R}$ be a complete, separated topological ring with a countable base of neighborhoods of zero consisting of open right ideals. Then every complex of flat left $\mathfrak{R}$-contramodules $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ can be included into a short exact sequence $0\longrightarrow\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow 0$ in $\mathsf{Com}(\mathfrak{R}{\operatorname{\mathsf{--Contra}}})$ with a contractible complex of flat cotorsion $\mathfrak{R}$-contramodules $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ and a complex of flat $\mathfrak{R}$-contramodules $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$.
@@
Let $\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be the direct sum of contractible two-term complexes $0\longrightarrow\overline{\mathfrak{C}}^{n}\overset{{\mathrm{id}}}{\longrightarrow}\overline{\mathfrak{C}}^{n}\longrightarrow 0$ with the cohomological grading defined by the rule that $\mathfrak{C}^{n}=\overline{\mathfrak{C}}^{n}\oplus\overline{\mathfrak{C}}^{n+1}$. Let $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be the morphism of complexes whose components are the monomorphisms $\mathfrak{F}^{n}\longrightarrow\overline{\mathfrak{C}}^{n}$ and the compositions $\mathfrak{F}^{n}\longrightarrow\mathfrak{F}^{n+1}\longrightarrow\overline{\mathfrak{C}}^{n+1}$. Then $\mathfrak{F}^{\text{\smaller\smaller$\scriptstyle\bullet$}}\longrightarrow\mathfrak{C}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ is a (termwise) monomorphism of complexes of $\mathfrak{R}$-contramodules. Let $\mathfrak{G}^{\text{\smaller\smaller$\scriptstyle\bullet$}}$ be the cokernel complex. Then there are short exact sequences of $\mathfrak{R}$-contramodules $0\longrightarrow\overline{\mathfrak{C}}^{n+1}\longrightarrow\mathfrak{G}^{n}\longrightarrow\overline{\mathfrak{G}}^{n}\longrightarrow 0$. Once again, Lemma 4.2 tells us that the $\mathfrak{R}$-contramodules $\mathfrak{G}^{n}$ are flat as extensions of flat $\mathfrak{R}$-contramodules. ∎